# floating point operation IVP initial value problem

### v1

@@ Abstract
Graph partitioning has many applications in power systems from decentralized state estimation to parallel simulation. Focusing on parallel simulation, optimal grid partitioning minimizes the idle time caused by different simulation times for the sub-networks and their components and reduces the overhead required to simulate the cuts. Partitioning a graph into two parts such that, for example, the cut is minimal and the sub-graphs have equal size is an NP-hard problem. In this paper we show how optimal partitioning of a graph can be obtained using quantum annealing (QA). We show how to map the requirements for optimal splitting to a quadratic unconstrained binary optimization (QUBO) formulation and test the proposed formulation using a current D-Wave QPU. We show that the necessity to find an embedding of the QUBO on current D-Wave QPUs limits the problem size to under 200 buses and notably affects the time-to-solution. We finally discuss the implications on near-term implementation of QA in combination to traditional CPU or GPU based simulation.
@@ I Introduction
With the transformation of the energy system towards renewable energies, the efforts required to model, simulate and control the system are steadily increasing. Whereas in the past a limited number of large power plants adjusted supply to demand, power balancing is increasingly characterized by a much larger number of smaller generation units, mostly based on fluctuating renewable energies, as well as more and more flexible consumers and prosumers. One way to manage the complexity of simulating and controlling these systems is to parallelize the problem by separating the overall system into smaller subsystems, which can then be solved on separate computing nodes. This network partitioning can also enable the use of flexible, distributed and adaptable smart grids monitoring [1] and control solutions [2]. Other grid partitioning applications include dividing into optimal balancing regions to reduce i. e. administrative and computational complexity and loop flows [3]; controlled islanding to prevent failure spreading [4] and diakoptics based state estimation [1]. In this paper we focus on the problem of partitioning a power system graph so to enable parallel simulation. While dependent on the specific simulation algorithm used, in general [5, 6, 7], an optimal partitioning for parallel simulation should ensure a good balancing of the computing loads and should minimize the overheads created by the partitioning.
@@
The problem of grid or graph partitioning, in two or more parts, is a fundamental problem in computer science and combinatorial optimization, known for its NP-hard complexity [8], which is hard to solve for classical computers and heuristics are usually applied [9]. Consequently, recent advances in quantum hardware have raised interest in exploring the potential of quantum computation as an alternative to established classical approaches. Thus, a new research field is emerging: Quantum computation for power systems [10, 11]. Recent advances in this field can be grouped into three categories:
@@
First, classical algorithms to solve systems of linear or nonlinear equations are replaced with quantum subroutines such as the HHL algorithm [12] or specially designed variational quantum circuits [13]. The HHL algorithm promises an exponential speed-up in the asymptotic limit if the problem is sparse [12] and the condition number can be bounded tightly [14]; however, it still faces input/output problems and requires circuits that are too deep for NISQ hardware. Nevertheless, proposed applications for HHL involve i. e. to solve DC and AC power flow equations [15, 16]. To circumvent some issues with the HHL algorithm, a variational circuit for a quantum electromagnetic transients program has been proposed [17]; however, quantum advantage has yet not been experimentally realized for any variational approach [13].
@@
Second, decision problems representing fundamental operational or planning problems in power systems are re-formulated to fit quantum optimization approaches. These involve the unit commitment problem [18], network reconfiguration in distribution grids [19] or optimal power flow [20]. However, on current hardware, only small problems can be tackled, which spiked recent interest in the decomposition of UC problem into smaller instances [21, 22].
@@
Third, quantum optimization is used to support classical simulation or optimization by, e.g., finding optimal cut in benders decomposition to solve large instances for the UC problem [23].
@@
The idea presented in this paper falls into the third category and explores graph partitioning to enhance the parallel simulation of power systems. The graph partitioning problem entails dividing the vertices of a graph into subgraphs while minimizing the number of edges between different subgraphs and/or achieving other specific objectives. Partitioning a graph into two parts, such that both subgraphs are of equal size, admits a natural quadratic unconstrained binary optimization (QUBO) formulation [24]. Thus, the problem fits established quantum optimization paradigms such as quantum annealing (QA) [25] or QAOA [26]. If advances in hardware are achievable, both approaches promise to solve large problem instances that are classically intractable. Partitioning into more than two parts based on the idea of network modularity can also be mapped to a QUBO; however, it requires costly integer encoding [27].
@@
The article is structured as follows: Section  II gives a short overview of QA. Section III reviews, combines and extends QUBO formulations for graph partitioning. Section IV gives a short overview of parallel simulation of power grids and its relation to graph partitioning, considering the specific algorithm of [5]. Section V combines the findings of the previous two sections to derive a tailored graph partitioning QUBO for parallel simulation. Section VI discusses the optimal and near-optimal solutions and current hardware limitations.
@@ II-A Quantum Annealing in a Nutshell
The idea of QA[25] is based on the quantum adiabatic theorem [28, 29] that roughly states that a closed system stays in its ground state if it undergoes an adiabatic transformation, that is a transformation that takes significantly longer than the time scales set by the internal dynamics. In quantum annealing, a system is initially prepared in a ground state of a mixer Hamiltonian $H_{M}$. Then the instantaneous Hamiltonian is ’slowly’ evolved towards the problem Hamiltonian $H_{P}$ in an annealing schedule given by
@@ II-C Time to (Best) Solution
As a consequence of sampling $N_{s}$ times, the annealing time $T_{A}$ is not the only parameter that affects the time to find a globally (or near) optimal solution, and there is a trade-off between longer annealing times and sampling more configurations. This trade-off is captured by the time to solution ($TTS$) [34, 35] which is defined as the ’time’ to reach a globally optimal (or best) solution, at least once with probability $p_{S}$ and is given by
@@
where $N_{opt}$ is the number of globally optimal (or: best) solutions in the sample set.
@@ III-A1 Minimal (Weighted) Cut
$\begin{split}Q_{cut}&=\sum_{n,m\in\mathcal{V}}z_{n}(-2A_{n,m}w_{n,m})z_{m}\\ &\,+\sum_{n\in\mathcal{V}}z_{n}\left(2\sum_{m\in\mathcal{V}}A_{n,m}w_{n,m}\right)z_{n},\end{split}$ (8)
@@ III-A3 Balancing the Sum of the Edge Weights
$\sum_{n\in\mathcal{V}}\underbrace{\left(\sum_{m\in\mathcal{V}}A_{n,m}w_{n,m}^{p}\right)}_{:=\alpha(n,p)}z_{n}=\sum_{n\in\mathcal{V}}\alpha(n,p)(1-z_{n}),$ (10)
where we have introduced the term $\alpha(n,p)$ to shorten the notation. Again, this constraint can be turned into a quadratic objective function $Q_{weights}(p)$ by the same steps as for $Q_{size}$. Note that the weights of the edges in the cut set are attributed to both parts and thus cancel each other out. The power $p$ can be used to tune the objective. If $p=1$, the sums of all weights in each part are balanced. For $p>>1$, the contributions of small weights are scaled down relatively to the maximal weights, such that balancing the maximal weights is emphasized.
@@ III-A4 Network Modularity
where $M$ is the number of edges, $deg(n)$ the degree of node $n$. To derive $Q_{mod}$ from [39] we have used that $\sum_{m}A_{n,m}=deg(n)$ and $\sum_{n}deg(n)=2M$.
@@ III-B QUBO using Linear Scalarization
The coefficients $\lambda_{i}$ have to be chosen a priori and effectively determine the importance of each objective as they set the energy scale. Different choices for the coefficients $\lambda_{i}$ can lead to different optimal points on the convex hull of the Pareto front [44].
@@ IV Parallel Simulation of Power Grids Based on network partitioning
$\begin{bmatrix}\begin{bmatrix}Y_{N1}&0\\ 0&Y_{N2}\end{bmatrix}&C\\ C^{T}&-Z_{L}\end{bmatrix}\begin{bmatrix}v_{N1}\\ v_{N2}\\ i_{L}\end{bmatrix}=\begin{bmatrix}h_{N1}\\ h_{N2}\\ -V_{L}\end{bmatrix},$ (13)
where $h_{Nk}$, $Y_{Nk}$, $v_{Nk}$ for $k\in\{1,2\}$ represent current injections, admittance matrix, nodal voltages (including currents chosen as independent variables as of MNA) in each sub-networks, respectively. $C$ represents the connectivity matrix of the link between partitions (referred to as ”cut” in the following), and $Z_{L}$, $i_{L}$, $V_{L}$ are the impedance matrix, current and voltage vectors over the cut, respectively. Reformulating Eq. (13) as
$\begin{split}\begin{bmatrix}1&0&Y_{N1}^{-1}C_{N1}\\ 0&1&Y_{N2}^{-1}C_{N2}\\ 0&0&C^{T}_{N1}C^{T}_{N2}+Z_{L}\\ \end{bmatrix}\begin{bmatrix}v_{N1}\\ v_{N2}\\ i_{L}\end{bmatrix}=\\ \begin{bmatrix}Y_{N1}^{-1}h_{N1}\\ Y_{N2}^{-1}h_{N2}\\ Y_{N1}h_{N1}+Y_{N2}h_{N2}+V_{L}\end{bmatrix},\end{split}$ (14)
shows that the solutions of each sub-network become independent of each other except for the cut. Hence, parallel computation of the sub-networks is possible after solving the equations over the cut. Additionally, finer parallelizations can be applied to each sub-network, such as the approaches in [6, 7, 47, 48]. Overall, each time step in simulating a partitioned network can be divided into three stages [5]:
@@
    minimizing the difference in computational cost among sub-networks to minimize the ’idle time’.
@@ V-A Estimation and Balancing of Computational Costs
We now need to estimate the computational costs of simulating the partitioned sub-networks to derive the objective function. For simplicity, we approximate these costs using the computation needed number of floating point operations.
@@ V-A1 Component costs
Dynamic components can be described as differential algebraic system of equations (DAE) and solved with numerical integration. Assuming explicit methods are used, we can discretize the integration as
$x_{k+1}$ $\leftarrow x_{k}+h\dot{x}_{k},$ (15)
$y$ $\leftarrow g(x),$ (16)
where Eq. (15) describes the integration of internal states $x,y$ of the component and Eq. (16) the injections from components to the grid. Therefore, we take the FLOPs in computing Eq. (15), Eq. (16) to represent the cost $c_{i}$ for solving the ith component.
A partitioning of the network into two sub-networks decomposes all components into three disjoint sets: the sets of all components in distinct parts (sub-network 1, 2), and in the cut, respectively. Let $C_{1(2)}$ and $C_{cut}$ be the index sets, e.g. the ith component is in part 1 if $i\in C_{1}$. Then the ’idle time’, i. e. mismatch of computing load in different sub-networks, is given by
@@ V-A2 Cut costs
$W_{cinj}=\sum_{i\in C_{cut}}c_{i}+4M_{c}(N-1)+M_{c},$
where $M_{c}$ is the number of edges in the cut.
@@ V-A3 Sub-network costs
Solving for the sub-networks is again done by LU decomposition. Therefore, the mismatch in solving sub-network is given by the difference between forward and backward substitution costs in each part:
$\begin{split}\Delta W_{networks}&=2(N_{1}^{2}-N_{2}^{2})-(N_{1}-N_{2})\\ &=2N\Delta N-\Delta N.\end{split}$ (19)

### v2

@@ Abstract
Graph partitioning has many applications in power systems from decentralized state estimation to parallel simulation. Focusing on parallel simulation, optimal grid partitioning minimizes the idle time caused by different simulation times for the sub-networks and their components and reduces the overhead required to simulate the cuts. Partitioning a graph into two parts such that, for example, the cut is minimal and the sub-graphs have equal size is an NP-hard problem. In this paper we show how optimal partitioning of a graph can be obtained using quantum annealing (QA). We show how to map the requirements for optimal splitting to a quadratic unconstrained binary optimization (QUBO) formulation and test the proposed formulation using a current D-Wave QPU. We show that the necessity to find an embedding of the QUBO on current D-Wave QPUs limits the problem size to under 200 buses and notably affects the time-to-solution. We finally discuss the implications of quantum hardware non-ideality on near term implementation in the simulation loop.
@@ I Introduction
The transition of energy systems towards renewable energies has significantly increased the complexity of modeling, simulating and controlling modern systems. Historically, power supply was adjusted to meet demand using a limited number of large power plants. However, in the future, power balancing is expected to be driven by a much larger number of smaller generation units, predominantly based on fluctuating renewable energy sources, alongside an increasing number of flexible consumers and prosumers in a grid with many more active grid elements. One approach to addressing the complexity of simulating and controlling these systems is to apply a parallelization strategy, which involves dividing the overall system into smaller subsystems that can be solved on separate computing nodes. This network partitioning can also facilitate the implementation of flexible, distributed and adaptive smart grid monitoring [1] and control solutions [2].
Other grid partitioning applications include dividing the power grid into optimally balanced regions to reduce i. e. administrative and computational complexity, and to simultaneously minimize loop flows [3]. Additionally, such partitioning can be applied to establish self-sufficient regions that allow for a certain level of electricity sharing [4], enable controlled islanding to prevent the spread of failures [5], and reduce the computational load in decentralized voltage control [6] and diakoptics-based state estimation [1].
The focus of this paper is the problem of partitioning a power system graph to enable parallel simulation. While dependent on the specific simulation algorithm employed, in general [7, 8, 9], an optimal partitioning for parallel simulation should ensure an adequate balancing of the computing loads and should minimize the overheads created by the partitioning.
@@
The problem of dividing a grid or graph into two or more parts is a fundamental problem in computer science and combinatorial optimization, known for its NP-hard complexity [10]. These problems are particularly difficult for classical computers to solve, which is why heuristics are usually applied [11]. Consequently, recent advancements in quantum hardware have raised interest in exploring the potential of quantum computation as an alternative to traditional classical approaches. This has led to the emergence of a new research field: Quantum computation for power systems [12, 13].
Recent advances in this field can be grouped into three categories:
@@
First, classical algorithms to solve systems of linear or nonlinear equations are being replaced with quantum subroutines such as the HHL algorithm [14] or specially designed variational quantum circuits [15]. The HHL algorithm offers the potential for exponential speed-up in the asymptotic limit, provided the problem is sparse [14] and the condition number can be tightly bounded [16]. However, it still faces challenges related to input/output processing and requires circuit depths that exceed the capabilities of current NISQ hardware. Despite these limitations, the HHL algorithm has been proposed for applications such as solving DC and AC power flow equations [17, 18]. To circumvent some issues with the HHL algorithm, a variational circuit for a quantum electromagnetic transients program has been proposed [19]. However, quantum advantage has yet not been experimentally demonstrated for any variational approach [15].
@@
Second, decision problems that represent fundamental operational or planning problems in power systems are being reformulated to fit quantum optimization approaches. These include the unit commitment (UC) problem [20], network reconfiguration in distribution grids [21], optimal power flow [22], and the aforementioned [4] grid partitioning into self-sufficient subgrids [23]. However, on current hardware, only small-scale problems can be tackled, which spiked recent interest in the decomposition of the UC problem into smaller instances [24, 25].
@@
Third, quantum optimization is being used to support classical simulation or optimization, for example, by finding optimal cuts in Benders decomposition to solve large instances of the UC problem [26].
@@
The idea presented in this paper falls into the third category and explores graph partitioning to enhance the parallel simulation of power systems. The graph partitioning problem involves dividing the vertices of a graph into subgraphs while minimizing the number of edges between different subgraphs and/or achieving other specific objectives. Partitioning a graph into two parts, such that both subgraphs are of equal size, naturally lends itself to a quadratic unconstrained binary optimization (QUBO) formulation [27]. Thus, the problem fits established quantum optimization paradigms such as quantum annealing (QA) [28] or QAOA [29]. If hardware advances are realized, both approaches promise to solve large problem instances that are classically intractable. Partitioning into more than two parts, based on the concept of network modularity, can also be mapped to a QUBO but it requires costly integer encoding [30]. This work focuses on partitioning a graph into two parts.
@@
The article is structured as follows: Section  II gives a short overview of QA. Section III reviews, combines and extends QUBO formulations for graph partitioning into two parts. Section IV gives a short overview of parallel simulation of power grids and its relation to graph partitioning, considering the specific algorithm of [7]. Section V combines the findings of the previous two sections to derive a QUBO for grid partitioning into two parts tailored to parallel simulation. We then briefly discuss possible extensions for partitioning into more than two parts. Section VI discusses the optimal and near-optimal solutions and current hardware limitations.
@@ II-A Quantum Annealing in a Nutshell
The idea of QA[28] is based on the quantum adiabatic theorem [31, 32] that roughly states that a closed system stays in its ground state if it undergoes an adiabatic transformation, that is a transformation that takes significantly longer than the time scales set by the internal dynamics. In quantum annealing, a system is initially prepared in a ground state of a mixer Hamiltonian $H_{M}$. Then the instantaneous Hamiltonian is “slowly” evolved towards the problem Hamiltonian $H_{P}$ in an annealing schedule given by
@@
For a more detailed introduction to QA and a discussion of industry applications, we refer to [37].
@@ II-C Time to (Best) Solution
As a consequence of sampling $N_{s}$ times, the annealing time $T_{A}$ is not the only parameter that affects the time to find a globally (or near) optimal solution, and there is a trade-off between longer annealing times and sampling more configurations. This trade-off is captured by the time to solution ($TTS$) [38, 39] which is defined as the “time” to reach a globally optimal (or best) solution, at least once with probability $p_{S}$ and is given by
@@
where $N_{opt}$ is the number of globally optimal (or: best) solutions in the sample set. In practice one assumes that the “success probability” of the heuristic solver, e.g. the quantum annealer, is below 1, that is, $N_{opt}<N_{s}$. Hence, Eq. 5 is well defined.
@@ III-A1 Minimal (Weighted) Cut
$\begin{split}Q_{cut}&=\sum_{n,m\in\mathcal{V}}z_{n}(-A_{n,m}w_{n,m})z_{m}\\ &\,+\sum_{n\in\mathcal{V}}z_{n}\left(\sum_{m\in\mathcal{V}}A_{n,m}w_{n,m}\right)z_{n},\end{split}$ (8)
@@ III-A3 Balancing the Sum of the Edge Weights
$\sum_{n\in\mathcal{V}}\underbrace{\left(\sum_{m\in\mathcal{V}}\frac{1}{2}A_{n,m}w_{n,m}^{p}\right)}_{:=\alpha(n,p)}z_{n}=\sum_{n\in\mathcal{V}}\alpha(n,p)(1-z_{n}),$
where we have introduced the term $\alpha(n,p)$ to shorten the notation. To avoid counting the edges within a sub-graph twice, we have multiplied both sides by $\frac{1}{2}$ and moved this factor into the definition of $\alpha(n,p)$. Note that the weights of the edges in the cut set are attributed to both parts and thus cancel each other out. The power $p$ can be used to tune the objective. If $p=1$, the sums of all weights in each part are balanced. For $p>>1$, the contributions of small weights are scaled down relatively to the maximal weights, such that balancing the maximal weights is emphasized.
Again, this constraint can be turned into a quadratic objective function $Q_{weights}(p)$ by the same steps as for $Q_{size}$. We get
$Q_{weights}(p)=\left(2\sum_{n\in\mathcal{V}}\alpha(n,p)z_{n}-\sum_{n\in\mathcal{V}}\alpha(n,p)\right)^{2}.$ (10)
@@ III-A4 Network Modularity
where $M$ is the number of edges, $deg(n)$ the degree of node $n$. To derive $Q_{mod}$ from [43] we have used that $\sum_{m}A_{n,m}=deg(n)$ and $\sum_{n}deg(n)=2M$ in order to eliminate the first order terms.
@@ III-B QUBO using Linear Scalarization
The coefficients $\lambda_{i}$ have to be chosen a priori and effectively determine the importance of each objective as they set the energy scale. Different choices for the coefficients $\lambda_{i}$ can lead to different optimal points on the convex hull of the Pareto front [48]. To solve the multi-objective optimization problem, one samples different Pareto optimal points using different sets of coefficients $\lambda_{i}$. The final solution is then selected by a “decision maker” using domain knowledge of the problem [49].
@@ IV Parallel Simulation of Power Grids Based on network partitioning
$\begin{bmatrix}\begin{bmatrix}Y_{N1}&0\\ 0&Y_{N2}\end{bmatrix}&C\\ C^{\top}&-Z_{L}\end{bmatrix}\begin{bmatrix}v_{N1}\\ v_{N2}\\ i_{L}\end{bmatrix}=\begin{bmatrix}h_{N1}\\ h_{N2}\\ -V_{L}\end{bmatrix},$ (13)
where $h_{Nk}$, $Y_{Nk}$, $v_{Nk}$ for $k\in\{1,2\}$ represent current injections, admittance matrix, nodal voltages (including currents chosen as independent variables as of MNA) in each sub-networks, respectively. $C$ represents the connectivity matrix of the link between partitions (referred to as“cut” in the following), and $Z_{L}$, $i_{L}$, $V_{L}$ are the impedance matrix, current and voltage vectors over the cut, respectively.
Consider the simplified system of equations
[…]
$\begin{bmatrix}1&Y^{-1}C\\ 0&-Z_{L}-C^{\top}Y^{-1}C\end{bmatrix}\begin{bmatrix}V\\ I_{L}\end{bmatrix}=\begin{bmatrix}Y^{-1}H\\ -V_{L}-C^{\top}Y^{-1}H\end{bmatrix}.$
Hence, Eq. (13) can be reformulated as
$\begin{split}\begin{bmatrix}1&0&Y_{N1}^{-1}C_{N1}\\ 0&1&Y_{N2}^{-1}C_{N2}\\ 0&0&C^{\top}_{N1}Y_{N1}^{-1}C_{N1}+C^{\top}_{N2}Y_{N2}^{-1}C_{N2}+Z_{L}\\ \end{bmatrix}\begin{bmatrix}v_{N1}\\ v_{N2}\\ i_{L}\end{bmatrix}&=\\ \begin{bmatrix}Y_{N1}^{-1}h_{N1}\\ Y_{N2}^{-1}h_{N2}\\ C_{N1}^{\top}Y^{-1}_{N1}h_{N1}+C_{N2}^{\top}Y^{-1}_{N2}h_{N2}+V_{L}\end{bmatrix}\,&\end{split}$ (14)
which shows that the solutions of each sub-network become independent of each other except for the cut. Hence, parallel computation of the sub-networks is possible after solving the equations over the cut. Additionally, finer parallelizations can be applied to each sub-network, such as the approaches in [8, 9, 52, 53]. Overall, each time step in simulating a partitioned network can be divided into three stages [7]:
@@
    minimizing the difference in computational cost among sub-networks to minimize the “idle time”.
@@ V-A Estimation and Balancing of Computational Costs
We now need to estimate the computational costs of simulating the partitioned sub-networks to derive the objective function. For simplicity, we approximate these costs using the number of computations (in the form of floating point operations) needed to evaluate the equations.
@@ V-A1 Component costs
The cost $c_{i}$ for solving the equations governing the $i$th component, e.g. a generator, depends on the numerical integration method used to evaluate the derivative or Jacobian of the model. Let $\sigma_{i,j}$ be the number of FLOPs for evaluating the equation governing the $j$th state variable in the $i$th component model. We have
$c_{i}=k\sum_{j}\sigma_{i,j}+\Delta_{i},$
where $k$ represents the number of times the derivatives or the Jacobian is evaluated. That is, $k$ depends on the number of stages or iterations of the integration method. Here, $\Delta_{i}$ quantifies the additional overhead such as vector addition, e.g. in $x\leftarrow x+dx$. For simplicity, we imply the use of Euler forward and ignore the additional overhead in this work, which yields
$c_{i}\approx\sum_{j}\sigma_{i,j}.$
We can illustrate the derivation of $c$ via a simple classical generator model [54] with two state variables
$\dot{\delta}$ $=\omega_{b}(\omega-\omega_{s}),$
$\dot{\omega}$ $=\frac{1}{2H}\left[(T_{m}-(\psi_{d}i_{q}-\psi_{q}i_{d})\right].$
The cost $\sigma_{\dot{\delta}}$ to evaluate the first state variable $\dot{\delta}$ can then be derived by counting the FLOPs, which here is $2$ for one subtraction and one multiplication. Similarly, $\sigma_{\dot{\omega}}$ is 5, since $\frac{1}{2H}$ is considered as one constant. Thus, the cost $c$ is $c=\sigma_{\dot{\delta}}+\sigma_{\dot{\omega}}=7$. The generator model implied in this work is a ninth-order model [54]. This model is significantly more complex, heuristically we evaluate the cost as $c=350$.
A partitioning of the network into two sub-networks decomposes all components into three disjoint sets: the sets of all components in distinct parts (sub-network 1, 2), and in the cut, respectively. Let $C_{1(2)}$ and $C_{cut}$ be the index sets, e.g. the ith component is in part 1 if $i\in C_{1}$. Then the “idle time”, i. e. mismatch of computing load in different sub-networks, is given by
@@ V-A2 Cut costs
$W_{cut,\,inj.}=\sum_{i\in C_{cut}}c_{i}+4M_{c}(N-1)+M_{c},$
where $M_{c}$ is the number of edges in the cut and $N$ is the total number of nodes/busses.
@@ V-A3 Sub-network costs
Solving for the sub-networks is again done by LU decomposition. Therefore, the mismatch in solving the sub-networks is given by the difference between forward and backward substitution costs in each part
$\begin{split}\Delta W_{net.}&=2(N_{1}^{2}-N_{2}^{2})-(N_{1}-N_{2})\\ &=2N\Delta N-\Delta N.\end{split}$ (17)
where $N_{1(2)}$ is the number of nodes/busses in sub-network $1$, or $2$ respectively.